# CogniBench: A Legal-inspired Framework and Dataset for Assessing Cognitive Faithfulness of Large Language Models

### v1

@@ 1 Introduction
  • In response to (C1), we introduce CogniBench (Section 2), the first knowledge-grounded dialogue dataset and framework for assessing cognitive faithfulness. Built through rigorous manual annotation, CogniBench provides sentence-level annotations with a legal-inspired assessment protocol to reduce subjectivity. To accommodate diverse application requirements, we define three increasingly rigorous faithfulness criteria: rational (plausible but unverifiable speculation), grounded (contextually supported evaluations), and conclusive (undeniable conclusions). The increasing levels of rigorousness cater to the diverse needs of various LLMs’ applications, from creative storytelling Zhou et al. (2023) to high-stakes medical diagnosis Singhal et al. (2025).
@@
Figure 3: (a)(b): Existing faithfulness assessment standards such as “Baseless” Niu et al. (2023) and “Subjective” Mishra et al. (2024) are ambiguous and insufficient for assessing cognitive statements. (c): We propose three increasingly rigorous assessment criteria (i.e. Rational, Grounded, Conclusive) to annotate cognitive statements. (d)(e): Human annotators assess each statement in (a) by making sequential decisions. Cognitive statements can be categorized as Misleading, Speculative, Reliable, or Conclusive based on what criteria are met. (See Section 2.2 for details)
@@ 2.2 A Legal-Inspired Tiered Framework for Assessing Cognitive Statements
Criterion 3. Conclusive: Whether a statement has no alternative interpretations could reasonably exist, free from subjective bias. This criterion is based on the legal requirement that “circumstantial evidence may be introduced, but the trier-of-fact must be satisfied beyond a reasonable doubt that the guilt of the accused is the only reasonable inference from the proven facts” v Villaroman (2016). Responses from AI systems such as AI Judger, AI Doctor, or AI Trader Yao et al. (2024); Singhal et al. (2025); Ding et al. (2024) should meet this standard. These systems make high-stakes decisions that directly impact human lives and livelihoods. Therefore, LLMs must make conclusive statements to minimize the risk of error or bias.
With these three criteria, a statement can be categorized into one of the following: Misleading, Speculative, Reliable, or Conclusive. For example, Cognitive Statement 1, labeled as Reliable (which might have been categorized as subjective or a hallucination under previous standards) is both Rational and Grounded, yet contains subjectivity. Despite this, it remains useful for AI assistant applications. Additional annotated examples are presented in Fig. 10.
@@
Our framework organizes the annotation process into a sequential decision-making structure (see full annotation protocol in Section A.5.2), guiding annotators through a sequential decision-making process to classify statements. The design ensures the annotator only considers a more rigorous criterion (e.g., grounded or conclusive) once previous less rigorous conditions (e.g., rational) are met, reducing cognitive load and the potential for conflicting interpretations. As shown in Table 1, this approach outperforms independent classification (i.e., direct ask annotator to categorize a statement into one type such as misleading, reliable) , achieving a 96.19% inter-annotator agreement (IAA) while also reducing QA instances by 48%.
@@ A.5.3 Full Instruction for Annotator
Rule 3: Conclusive: Whether the statement is the only reasonable conclusion supported by indisputable evidence, free from subjective bias.
@@ A.5.4 Examples of Annotation
We provide annotated examples in Fig. 10 for four types of cognitive statements: Misleading, Speculative, Reliable, and Conclusive. The context has been excerpted for ease of reading.

### v2

@@ 1 Introduction
  • In response to (C1), we introduce CogniBench (Section 2), the first knowledge-grounded dialogue dataset and framework for assessing cognitive faithfulness. Built through rigorous manual annotation, CogniBench provides sentence-level annotations with a legal-inspired assessment protocol to reduce subjectivity. To accommodate diverse application requirements, we define three increasingly rigorous faithfulness criteria: rational (plausible but unverifiable speculation), grounded (contextually supported evaluations), and unequivocal (undeniable conclusions). The increasing levels of rigorousness cater to the diverse needs of various LLMs’ applications, from creative storytelling Zhou et al. (2023) to high-stakes medical diagnosis Singhal et al. (2025).
@@
Figure 3: (a)(b): Existing faithfulness assessment standards such as “Baseless” Niu et al. (2023) and “Subjective” Mishra et al. (2024) are ambiguous and insufficient for assessing cognitive statements. (c): We propose three increasingly rigorous assessment criteria (i.e. Rational, Grounded, Unequivocal) to annotate cognitive statements. (d)(e): Human annotators assess each statement in (a) by making sequential decisions. Cognitive statements can be categorized as Misleading, Speculative, Reliable, or Unequivocal based on what criteria are met. (See Section 2.2 for details)
@@ 2.2 A Legal-Inspired Tiered Framework for Assessing Cognitive Statements
Criterion 3. Unequivocal: Whether a statement has no alternative interpretations could reasonably exist, free from subjective bias. This criterion is based on the legal requirement that “circumstantial evidence may be introduced, but the trier-of-fact must be satisfied beyond a reasonable doubt that the guilt of the accused is the only reasonable inference from the proven facts” v Villaroman (2016). Responses from AI systems such as AI Judger, AI Doctor, or AI Trader Yao et al. (2024); Singhal et al. (2025); Ding et al. (2024) should meet this standard. These systems make high-stakes decisions that directly impact human lives and livelihoods. Therefore, LLMs must make unequivocal statements to minimize the risk of error or bias.
With these three criteria, a statement can be categorized into one of the following: Misleading, Speculative, Reliable, or Unequivocal. For example, Cognitive Statement 1, labeled as Reliable (which might have been categorized as subjective or a hallucination under previous standards) is both Rational and Grounded, yet contains subjectivity. Despite this, it remains useful for AI assistant applications. Additional annotated examples are presented in Fig. 10.
@@
Our framework organizes the annotation process into a sequential decision-making structure (see full annotation protocol in Section A.5.2), guiding annotators through a sequential decision-making process to classify statements. The design ensures the annotator only considers a more rigorous criterion (e.g., grounded or unequivocal) once previous less rigorous conditions (e.g., rational) are met, reducing cognitive load and the potential for conflicting interpretations. As shown in Table 1, this approach outperforms independent classification (i.e., direct ask annotator to categorize a statement into one type such as misleading, reliable) , achieving a 96.19% inter-annotator agreement (IAA) while also reducing QA instances by 48%.
@@ A.5.3 Full Instruction for Annotator
Rule 3: Unequivocal: Whether the statement is the only reasonable conclusion supported by indisputable evidence, free from subjective bias.
@@ A.5.4 Examples of Annotation
We provide annotated examples in Fig. 10 for four types of cognitive statements: Misleading, Speculative, Reliable, and Unequivocal. The context has been excerpted for ease of reading.